### v1

@@ 3 Several propositions
The subword $u^{\prime}$ is not necessarily unique, since we can take some 0’s from blocks in a different way. For example, if $u^{\prime}$ has only one zero from a block $0^{s}$, then there are $s$ ways to take this zero.
@@
Note that either $v$ does not have a subword which equal to $u^{\prime}$ or there is a subword $v^{\prime}=u^{\prime}$ of $v$. In the first case $u^{\prime}$ is the subword of the word $u$ and is not a subword of $v$ and $|u^{\prime}|\leqslant\frac{3}{4}n+3$. In the second case we can overlay by 1’s the word $v$ on $u$ such that such that $v^{\prime}$ and $u^{\prime}$ coincide. Since $v\neq u$, then there is a block $0^{s}$ in $u$ (probably empty) which is smaller than the corresponding block of 0’s in $v$ ($s\leqslant y$). We then add a block $0^{s+1}$ to $v^{\prime}$ to the corresponding place. We get a subword $v^{\prime\prime}$ of the word $v$ which is not a subword of $u$, and $|v^{\prime\prime}|\leqslant|v^{\prime}|+s+1\leqslant|v^{\prime}|+y+1\leqslant\frac{3}{4}n+4$.
@@
In this section, we provide a proof of Theorem 1.1. In Subsection 4.1 we provide a general structure of the proof, splitting it into several lemmas, and give some auxiliary notation used throughout the proof. In Subsection 4.2, we state and prove lemmas constituting the proof.
@@ 4.1 Notation and general structure of the proof
In this subsection, we fix some notation and give a general description of the proof of the main results of this paper, Theorem 1.1. Namely, let $u\neq v$ be two cyclic words of length $n$. We will prove that there is a word $w$ of length at most $\frac{3}{4}n+4$ such that $w$ is a subword of only one of the words $u$ and $v$. By Corollary 3.9, it remains to prove the theorem when $n_{0,u}=n_{0,v}=n_{0}$, $n_{1,u}=n_{1,v}=n_{1}$ and $l_{u}=l_{v}=l$. Without loss of generality we can assume that $n_{0}\geqslant n_{1}$ (or, equivalently, $n_{1}\leqslant\frac{n}{2}$).
Recall that $0^{x_{u}}$ and $0^{x_{v}}$ are the longest blocks of 0’s in $u$ and $v$, respectively. Let $y=y_{u}$ be the length of the longest block of 00’s in $u$ which is shorter than $0^{x_{u}}$ (if all the blocks have the same length, we set $y=0$). We let $a$ and $b$ denote the numbers of blocks $0^{x_{u}}$ and $0^{x_{v}}$ in $u$ and $v$, respectively. We will divide the proof of the theorem into five lemmas according to different cases as follows:
@@
  • $u_{long}$ is a non-special and non-periodic word and $a\geqslant 3$, or, symmetrically, $v_{long}$ is non-special and non-periodic word, $b\geqslant 3$. (Lemma 4.2).
  • $u_{long}$ is periodic and non-special, $a\geqslant 3$, or, symmetrically, $v_{long}$ is periodic and non-special, $b\geqslant 3$ (Lemma 4.4),
  • If one of the following two conditions holds (Lemma 4.5):
    $\bullet$ both words are special and $a=b=2$.
    $\bullet$ one of the words $u,v$ is a non-special word with at most two blocks $0^{x}$, and the other word is either special or contains at most two blocks $0^{x}$.
  • $u_{long}$ and $v_{long}$ are special words (Lemma 4.7).
It is not hard to see that all the cases are covered. In the first lemma we prove the theorem for words which have different sizes of big blocks. In the second and third lemmas we prove the theorem in the case when one of the words non-special and has at least three big blocks. In fourth and fifth lemmas we prove theorem in the case when either both words are special or one of the word non-special and has one or two big blocks.
@@ 4.2 Lemmas constituting the proof
Let $x_{u}\neq x_{v}$. Then there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
Proof. We will consider two cases:
1. $a\geqslant 2$ or $b\geqslant 2$.
Without loss of generality assume that $a\geqslant 2$. If $x_{u}>x_{v}$, then consider the word $u_{1}=1^{n_{1}}0^{x_{u}}$. This word is the subword of the word $u$ and is not a subword of $u$. Also $|u_{1}|=n_{1}+x_{u}$. Since $a\geqslant 2$, then $x_{u}\leqslant\frac{ax_{u}}{2}\leqslant\frac{n_{0}}{2}$. So, $|u_{1}|=n_{1}+x_{u}\leqslant n_{1}+\frac{n_{0}}{2}=\frac{n}{2}+\frac{n_{1}}{2}\leqslant\frac{3}{4}n$. If $x_{u}<x_{v}$, then consider the word $u_{1}=1^{n_{1}}0^{x_{u}+1}$. This word is the subword of the word $v$ and is not a subword of $u$, and its length is $|u_{1}|=n_{1}+x_{u}+1\leqslant n_{1}+\frac{n_{0}}{2}+1\leqslant\frac{3}{4}n+1$. In both cases the lemma is proved.
2. $a=b=1$.
Without loss of generality we assume that $x_{u}>x_{v}$. The words $u_{1}=1^{n_{1}}0^{x_{v}+1}$ and $u_{2}=(01)^{l-1}0^{x_{v}+1}1$ are subwords of the word $u$ and are not subwords of $v$. If $|u_{1}|\leqslant\frac{3}{4}n+4$ or $|u_{2}|\leqslant\frac{3}{4}n+4$, then we have a required subword. Otherwise $|u_{1}|=n_{1}+x_{v}+1>\frac{3}{4}n+4$ and $|u_{2}|=2l+x_{v}>\frac{3}{4}n+4$. If $x_{v}\leqslant\frac{n_{0}}{2}$, then $|u_{1}|\leqslant n_{1}+\frac{n_{0}}{2}+1\leqslant\frac{3}{4}n+1$. So $x_{u}>x_{v}>\frac{n_{0}}{2}$ and each block of 0’s in $u$ except for $0^{x_{u}}$ contains less than $x_{v}$ occurrences of 0 (in particular the block $0^{y}$).
We now overlay by 1’s the word $v$ on $u$ such that the blocks $0^{x_{u}}$ and $0^{x_{v}}$ coincide. Since $x_{u}>x_{v}$, then for this overlay there are two neighboring 1’s such that $v$ has more 0’s between them than $u$ has. We let $t$ and $s>t$ denote the numbers of occurrences of 0’s between them in $u$ and $v$, respectively. We let $p$ denote the number of occurrences of 1’s between this block of 0’s and the block $0^{x_{v}}$ in $v$. Consider the word $v_{1}=1^{p}0^{y+1}1^{n_{1}-p}0^{t+1}$. The word $v_{1}$ is a subword of the word $v$ and is not a subword of $u$. Since $u$ has at least $l+(x_{u}-1)+(y-1)$ occurrences of 0 (that is, $n_{0}\geqslant l+(x_{u}-1)+(y-1)$) and $y\geqslant t$, then
$|u_{1}|+|u_{2}|+|u_{3}|=(n_{1}+x_{v}+1)+(2l+x_{v})+(n_{1}+y+1+t+1)\leqslant$
@@
So, the length of at least one of the words $u_{1}$, $u_{2}$, $v_{1}$ at most $\frac{2n+7}{3}<\frac{3}{4}n+4$.
Lemma is proved. ∎
In the following text we will assume that $x_{u}=x_{v}=x$. Note that since $n_{0,u}=n_{0,v}=n_{0}$, $n_{1,u}=n_{1,v}=n_{1}\leqslant\frac{n}{2}$ and $l_{u}=l_{v}$, then $x\neq 1$, otherwise both words are equal to $(01)^{k}$. So, we will also assume that $x\geqslant 2$. Recall that $u_{long}$ and $v_{long}$ are subwords of $u$ and $v$, respectively, which contain all 1’s and all big blocks $0^{x}$.
@@
Let $u_{long}$ be a non-special and non-periodic word and let $a\geqslant 3$. Then there exists a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
Proof. Due to Corollary 3.5, there exists a unique subword $u_{fix}$ of $u$ such that $u_{fix}$ has $n_{1}$ occurrences of 1 and at most $\frac{a}{2}+1$ blocks $0^{x}$. The proof of the lemma is split into the following cases:
1. $n_{1}+\frac{ax}{2}+2x\leqslant\frac{3}{4}n+3$, $a\geqslant 4$.
2. $n_{1}+\frac{ax}{2}+2x\geqslant\frac{3}{4}n+3$, $a\geqslant 4$
2.1. $a\geqslant 5$.
2.1.1. $u_{zeros}$ is not periodic
2.1.2. $u_{zeros}$ is periodic
2.2. $a=4$
3. $a=3$
@@
1. $n_{1}+\frac{ax}{2}+x+y\leqslant\frac{3}{4}n+3$, $a\geqslant 4$
The subword $u_{fix}$ of $u$ is unique and $|u_{fix}|+y+1\leqslant n_{1}+(\frac{a}{2}+1)x+y+1\leqslant n_{1}+\frac{ax}{2}+x+y+1\leqslant\frac{3}{4}n+4$. Then applying Proposition 3.11 to the word $u_{fix}$ we get that there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
2. $n_{1}+\frac{ax}{2}+x+y\geqslant\frac{3}{4}n+3$, $a\geqslant 4$
2.1. $a\geqslant 5$
Recall that $n_{long}$ and $n_{short}$ are the numbers of 0’s in big blocks (blocks $0^{x}$) and small blocks (blocks $0^{<x}$), respectively. Notice that $n_{0}=n_{long}+n_{short}=ax+n_{short}$, $n=n_{1}+n_{0}=n_{1}+ax+n_{short}.$ So,
$n_{short}=n-n_{1}-ax.$ (1)
The word $u$ has $a$ blocks $0^{x}$ and at most $n_{short}-y+1$ small blocks. So,
$l\leqslant n_{short}-y+1+a.$ (2)
Since $n_{1}+\frac{ax}{2}+x+y\geqslant\frac{3}{4}n+3$, then (1) and (2) imply that
$l\leqslant n_{short}-y+1+a=n-n_{1}-ax-y+a+1=n-(n_{1}+\frac{ax}{2}+x+y)-\frac{ax}{2}+x+a+1\leqslant$
$\ \leqslant\frac{n}{4}-\frac{ax-2x-2a+4}{2}=\frac{n}{4}-\frac{(a-2)(x-2)}{2}.$ (3)
Consider the words $u_{zeros}$ and $v_{zeros}$ (recall that they are subwords of $u$ and $v$ which contain all 1’s and one zero from each block of 0’s). Notice that $|u_{zeros}|=n_{1}+l\leqslant\frac{n}{2}+\frac{n}{4}-\frac{(a-2)(x-2)}{2}\leqslant\frac{3}{4}n$, since $a\geqslant 2,x\geqslant 2$. If $v_{zeros}\neq u_{zeros}$, then $v_{zeros}$ is the is a subword of the word $v$ and is not a subword of $u$, and $|v_{zeros}|\leqslant\frac{3}{4}n$. So it remains to prove the lemma in the case $v_{zeros}=u_{zeros}$. Consider two cases.
2.1.1. $u_{zeros}$ is not periodic.
In this case there is only one way to take $u_{zeros}$ in $u$ accurate to the selection of one zero from each block of 0’s. Since $n_{1}\leqslant\frac{n}{2}$ and by (3), we have
$|u_{zeros}|+y+1=n_{1}+l+y+1<n_{1}+\frac{n}{4}-\frac{ax-2x-2a+4}{2}+x+1\leqslant$
@@ 2.1. $a\geqslant 5$
since $a\geqslant 5$ and $x\geqslant 2$. Applying Proposition 3.11 to the word $u_{zeros}$ we get that there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
2.1.2. $u_{zeros}$ is periodic
The word $u_{zeros}$ is of the following form: $u_{zeros}=(1^{\alpha_{1}}01^{\alpha_{2}}0\cdots 1^{\alpha_{t}}0)^{\frac{l}{t}}$, where $\frac{l}{t}\geqslant 2$ is an integer and $t$ is minimal. Consider the word $u_{zeros}^{\prime}=(1^{\alpha_{1}}01^{\alpha_{2}}0\cdots 1^{\alpha_{t}}0)(10)^{l-t}$. The word $01^{\alpha_{1}}01^{\alpha_{2}}0\cdots 1^{\alpha_{t}}$ 0 contains $\frac{n_{1}}{\frac{l}{t}}$ occurrences of 1, hence $u_{zeros}^{\prime}$ contains $\frac{n_{1}\cdot t}{l}+l-t$ occurrences of 1 and $l$ occurrences of 0. The function $f(t)=\frac{n_{1}\cdot t}{l}+l-t$ is increasing and $t\leqslant\frac{l}{2}$ (since $\frac{l}{t}\geqslant 2$ is integer), so $u_{zeros}^{\prime}$ contains at most $\frac{n_{1}+l}{2}$ occurrences of 1. We add to $u_{zeros}^{\prime}$ all blocks $0^{x}$ from $u_{fix}$ defined in the beginning of the proof. Let $u_{zeros}^{\prime\prime}$ denote the obtained subword of $u$. There is only one way to take $u_{zeros}^{\prime\prime}$ in $u$ accurate to the selection of 1’s and 0’s from blocks in which we get not all symbols. Since $l\leqslant n_{short}-y+a+1$, then $n=n_{1}+ax+n_{short}\geqslant n_{1}+ax+l+y-a-1$. From these inequalities and (3) it follows that
$|u_{zeros}^{\prime\prime}|+y+1\leqslant\frac{n_{1}+l}{2}+l+\left(\frac{a+2}{2}\right)(x-1)+y+1=$
$=\frac{n_{1}+ax+l+y-a-1}{2}+l+x+\frac{y}{2}+\frac{1}{2}<\frac{n}{2}+\frac{n}{4}-\frac{ax-2x-2a+4}{2}+\frac{3x}{2}=$
@@ 2.1.2. $u_{zeros}$ is periodic
since $a\geqslant 5$ and $x\geqslant 2$. Applying Proposition 3.11 to the word $u_{zeros}^{\prime\prime}$, we get that there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
2.2. $a=4$
In this case the proof is similar to the proof in case 2.1. All inequalities from case 2.1 hold true, except for the inequality (5) in the case 2.1.2. If we prove that $l\leqslant\frac{n}{4}-\frac{ax-2x-2a+4}{2}-\frac{y}{2}+1$, then we can rewrite inequality (5) as
$|u_{zeros}^{\prime\prime}|+y+1\leqslant\frac{n_{1}+ax+l+y-a-1}{2}+l+x+\frac{y}{2}+\frac{1}{2}<\frac{n}{2}+\frac{n}{4}-\frac{ax-2x-2a+4}{2}-$
$-\frac{y}{2}+1+x+\frac{x}{2}=\frac{3}{4}n-\frac{ax-2a-4x+2}{2}=\frac{3}{4}n-\frac{(a-4)(x-2)}{2}+3=\frac{3}{4}n+3,$
which gives us the proof in this similarly to the proof in case 2.1.
@@ 2.2. $a=4$
$l\leqslant\frac{n}{4}-x+3-\frac{y}{2}$ (3’)
We now consider several cases according to the number of blocks of 0’s of length at least $\frac{y}{2}$ in $u$ and in $v$. Since in case 2.2. we are proving now we have $a=4$, i.e., the number of blocks of 0’s of length $x$ (the longest blocks) in $u$ is $4$, $u_{long}$ is non-special by the conditions of the lemma, we have that the number of blocks of length at least $\frac{y}{2}$ in $u$ is at least $5$.
If there are at least six blocks in $u$ with lengths at least $0^{\frac{y}{2}}$, i.e. four blocks $0^{x}$, one block $0^{y}$ and at least one block $0^{\frac{y}{2}}$, then we can rewrite inequality (3) as $l\leqslant n_{short}-y-(\frac{y}{2}-1)+1+a$. So, in this case we have $l\leqslant\frac{n}{4}-\frac{ax-2x-2a+4}{2}-\frac{y}{2}+1$.

### v2

@@ 3 Preliminary observations and auxiliary statements
Note that in the second case, when $u^{\prime}$ is a subword of $u$ which contains $n_{1}$ occurrences of 1 and such that the only turn translating $u^{\prime}$ to an equal subword is the identity map, $u^{\prime}$ does not have to be unioccurent, since we can choose different 0’s from blocks of length greater than $1$.
@@
Note that either $v$ does not have a subword which is equal to $u^{\prime}$ or there is a subword $v^{\prime}=u^{\prime}$ of $v$. In the first case $u^{\prime}$ is a subword of the word $u$ and is not a subword of $v$ and $|u^{\prime}|\leqslant\frac{3}{4}n+3$. In the second case we can consider a 1-overlay of $v$ on $u$ such that $v^{\prime}$ and $u^{\prime}$ coincide. Since $v\neq u$, there is a block $0^{s}$ in $u$ ($s$ is the length of the block, and we set $s=0$ if the block is empty) which is smaller than the corresponding block of 0’s in $v$ (we let $y$ denote the length of this block, so that $s\leqslant y$). We add the block $0^{s+1}$ to $v^{\prime}$ to the corresponding place. We get a subword $v^{\prime\prime}$ of the word $v$ which is not a subword of $u$, and $|v^{\prime\prime}|\leqslant|v^{\prime}|+s+1\leqslant|v^{\prime}|+y+1\leqslant\frac{3}{4}n+4$. ∎
@@
In this section, we provide a proof of Theorem 1.1. In Subsection 4.1 we give a general structure of the proof, splitting it into several lemmas, and introduce some auxiliary notation used throughout the proof. In Subsection 4.2, we state and prove lemmas constituting the proof.
@@ 4.1 Notation and general structure of the proof
In this subsection, we fix some notation and give a general description of the proof of the main result of this paper, Theorem 1.1. Namely, let $u\neq v$ be two cyclic words of length $n$. We will prove that there is a word $w$ of length at most $\frac{3}{4}n+4$ such that $w$ is a distinguishing subword for the words $u$ and $v$. By Corollary 3.5, it remains to prove the theorem when $n_{0,u}=n_{0,v}=n_{0}$, $n_{1,u}=n_{1,v}=n_{1}$ and $l_{u}=l_{v}=l$. Without loss of generality we can assume that $n_{0}\geqslant n_{1}$ (or, equivalently, $n_{1}\leqslant\frac{n}{2}$).
Recall that $0^{x_{u}}$ and $0^{x_{v}}$ are the longest blocks of 0’s in $u$ and $v$, respectively, and that $y=y_{u}$ is the length of the longest block of 0’s in $u$ which is shorter than $0^{x_{u}}$ (if all the blocks have the same length, we set $y=0$). We let $a$ and $b$ denote the numbers of blocks $0^{x_{u}}$ and $0^{x_{v}}$ in $u$ and $v$, respectively. We divide the proof of our theorem into five lemmas according to different cases as follows:
@@
  • $u_{\operatorname{long}}$ is a neither special nor periodic word with $a\geqslant 3$, or, analogously, $v_{\operatorname{long}}$ is a neither special nor periodic word with $b\geqslant 3$ (Lemma 4.2).
  • $u_{\operatorname{long}}$ is periodic and not special, $a\geqslant 3$, or, analogously, $v_{\operatorname{long}}$ is periodic and not special, $b\geqslant 3$ (Lemma 4.4).
  • one of the words $u_{\operatorname{long}},v_{\operatorname{long}}$ is not a special word with at most two blocks $0^{x}$, and the other word is either special or contains at most two blocks $0^{x}$ (Lemma 4.5).
  • $u_{\operatorname{long}}$ and $v_{\operatorname{long}}$ are special words (Lemma 4.6).
It is not hard to see that all the cases are covered. In the first lemma we prove the theorem for words which have different sizes of big blocks. In the second and third lemmas we prove the theorem in the case when one of the words is not special and has at least three big blocks. In the fourth and fifth lemmas we prove the theorem in the case when either both words are special or one of the word is not special and has one or two big blocks.
@@ 4.2 Lemmas constituting the proof
Let $x_{u}\neq x_{v}$. Then $u$ and $v$ have a distinguishing subword of length at most $\frac{3}{4}n+4$.
Proof. We consider two cases:
Case 1. $a\geqslant 2$ or $b\geqslant 2$.
Without loss of generality assume that $a\geqslant 2$. If $x_{u}>x_{v}$, then consider the word $u_{1}=1^{n_{1}}0^{x_{u}}$. This word is the subword of the word $u$ and is not a subword of $u$. Since $a\geqslant 2$, we have $x_{u}\leqslant\frac{ax_{u}}{2}\leqslant\frac{n_{0}}{2}$. So, $|u_{1}|=n_{1}+x_{u}\leqslant n_{1}+\frac{n_{0}}{2}=\frac{n}{2}+\frac{n_{1}}{2}\leqslant\frac{3}{4}n$. If $x_{u}<x_{v}$, then consider the word $u_{1}=1^{n_{1}}0^{x_{u}+1}$. This word is a subword of the word $v$ and is not a subword of $u$, and its length is $|u_{1}|=n_{1}+x_{u}+1\leqslant n_{1}+\frac{n_{0}}{2}+1\leqslant\frac{3}{4}n+1$. In both cases we have a desired distinguishing subword.
Case 2. $a=b=1$.
Without loss of generality we assume that $x_{u}>x_{v}$. The words $u_{1}=1^{n_{1}}0^{x_{v}+1}$ and $u_{2}=(01)^{l-1}0^{x_{v}+1}1$ are subwords of the word $u$ and are not subwords of $v$. If $|u_{1}|\leqslant\frac{3}{4}n+4$ or $|u_{2}|\leqslant\frac{3}{4}n+4$, then we have a required subword. Otherwise $|u_{1}|=n_{1}+x_{v}+1>\frac{3}{4}n+4$ and $|u_{2}|=2l+x_{v}>\frac{3}{4}n+4$. If $x_{v}\leqslant\frac{n_{0}}{2}$, then $|u_{1}|\leqslant n_{1}+\frac{n_{0}}{2}+1\leqslant\frac{3}{4}n+1$. So, it remains to consider the case $x_{u}>x_{v}>\frac{n_{0}}{2}$ and each block of 0’s in $u$ except for $0^{x_{u}}$ contains less than $x_{v}$ occurrences of 0 (in particular, the block $0^{y}$).
We now consider a 1-overlay of $v$ on $u$ such that the blocks $0^{x_{u}}$ and $0^{x_{v}}$ coincide. Since $x_{u}>x_{v}$, there are two neighboring 1’s such that $v$ has more 0’s between them than $u$ has for this overlay. We let $t$ and $s>t$ denote the numbers of occurrences of 0 between them in $u$ and $v$, respectively. We let $p$ denote the number of occurrences of 1 between this block of 0’s and the block $0^{x_{v}}$ in $v$. Consider the word $v_{1}=1^{p}0^{y+1}1^{n_{1}-p}0^{t+1}$. The word $v_{1}$ is a subword of the word $v$ and is not a subword of $u$. Since $u$ has at least $l+(x_{u}-1)+(y-1)$ occurrences of 0 (that is, $n_{0}\geqslant l+(x_{u}-1)+(y-1)$) and $y\geqslant t$, we have
$|u_{1}|+|u_{2}|+|v_{1}|=(n_{1}+x_{v}+1)+(2l+x_{v})+(n_{1}+y+1+t+1)\leqslant$
@@
So, the length of at least one of the words $u_{1}$, $u_{2}$, $v_{1}$ is at most $\frac{2n+7}{3}<\frac{3}{4}n+4$. ∎
In the following text we assume that $x_{u}=x_{v}=x$. Note that since $n_{0,u}=n_{0,v}=n_{0}$, $n_{1,u}=n_{1,v}=n_{1}\leqslant\frac{n}{2}$ and $l_{u}=l_{v}$, we have $x\neq 1$. Otherwise both words are equal to $(01)^{k}$. So, we also assume that $x\geqslant 2$. Recall that $u_{\operatorname{long}}$ and $v_{\operatorname{long}}$ are subwords of $u$ and $v$, respectively, which contain all 1’s and all big blocks $0^{x}$.
@@
Let $u_{\operatorname{long}}$ be a neither special nor periodic word with $a\geqslant 3$. Then there exists a distinguishing subword of length at most $\frac{3}{4}n+4$ for $u$ and $v$.
Proof. Due to Corollary 3.9, there exists a unioccurrent subword $u_{\operatorname{fix}}$ of $u_{\operatorname{long}}$ such that $u_{\operatorname{fix}}$ has $n_{1}$ occurrences of 1 and at most $\frac{a}{2}+1$ blocks $0^{x}$. It is easy to see that $u_{\operatorname{fix}}$ is also a unioccurrent subword of $u$. The proof of the lemma is split into several cases as follows. Cases 1 and 2 correspond to $a\geqslant 4$; Case 1 gives a proof under the condition $n_{1}+\frac{ax}{2}+2x<\frac{3}{4}n+3$, and Case 2 treats the opposite inequality. Case 2 is divided into subcases 2.1 and 2.2 corresponding to $a\geqslant 5$ and $a=4$, respectively. Case 2.1 is further subdivided to subcases 2.1.1 and 2.1.2 depending on whether $u_{\operatorname{zeros}}$ is periodic or not. Case 3 corresponds to $a=3$.
@@
Case 1. $n_{1}+\frac{ax}{2}+x+y<\frac{3}{4}n+3$, $a\geqslant 4$.
The subword $u_{\operatorname{fix}}$ of $u$ is unioccurrent and $|u_{\operatorname{fix}}|+y+1\leqslant n_{1}+(\frac{a}{2}+1)x+y+1\leqslant n_{1}+\frac{ax}{2}+x+y+1<\frac{3}{4}n+4$. Then applying Proposition 3.11 to the word $u_{\operatorname{fix}}$ we get that there is a distinguishing subword of length at most $\frac{3}{4}n+4$ for $u$ and $v$.
Case 2. $n_{1}+\frac{ax}{2}+x+y\geqslant\frac{3}{4}n+3$, $a\geqslant 4$.
Case 2.1. $a\geqslant 5$.
Recall that $n_{\operatorname{long}}$ and $n_{\operatorname{short}}$ are the numbers of 0’s in big blocks (blocks $0^{x}$) and small blocks (blocks $0^{<x}$), respectively. Notice that $_{0}=n_{\operatorname{long}}+n_{\operatorname{short}}=ax+n_{\operatorname{short}}, n=n$$n_{1}+n_{0}=n_{1}+ax+n_{\operatorname{short}}$. So,
$n_{\operatorname{short}}=n-n_{1}-ax.$ (1)
The word $u$ has $a$ blocks $0^{x}$ and at most $n_{\operatorname{short}}-y+1$ small blocks. So,
$l\leqslant n_{\operatorname{short}}-y+1+a.$ (2)
Since $n_{1}+\frac{ax}{2}+x+y\geqslant\frac{3}{4}n+3$, (1) and (2) imply that
$l\leqslant n_{\operatorname{short}}-y+1+a=n-n_{1}-ax-y+a+1=n-(n_{1}+\frac{ax}{2}+x+y)-\frac{ax}{2}+x+a+1\leqslant$
[…]
Consider the words $u_{\operatorname{zeros}}$ and $v_{\operatorname{zeros}}$ (recall that they are subwords of $u$ and $v$ which contain all 1’s and one zero from each block of 0’s). Notice that $|u_{\operatorname{zeros}}|=n_{1}+l\leqslant\frac{n}{2}+\frac{n}{4}-\frac{(a-2)(x-2)}{2}\leqslant\frac{3}{4}n$, since $a\geqslant 2,x\geqslant 2$. If $v_{\operatorname{zeros}}\neq u_{\operatorname{zeros}}$, then $u_{\operatorname{zeros}}$ is a subword of the word $u$ and is not a subword of $v$, and $|u_{\operatorname{zeros}}|\leqslant\frac{3}{4}n$. So it remains to prove the lemma in the case $v_{\operatorname{zeros}}=u_{\operatorname{zeros}}$. Consider two cases.
Case 2.1.1. $u_{\operatorname{zeros}}$ is not periodic.
In this case there is only one way to take an occurrence of $u_{\operatorname{zeros}}$ in $u$ modulo the selection of one zero from each block of 0’s. Since $n_{1}\leqslant\frac{n}{2}$ and by (3), we have
$|u_{\operatorname{zeros}}|+y+1=n_{1}+l+y+1<n_{1}+\frac{n}{4}-\frac{ax-2x-2a+4}{2}+x+1\leqslant$
@@
since $a\geqslant 5$ and $x\geqslant 2$. Applying Proposition 3.11 to the word $u_{\operatorname{zeros}}$ we get that there is a distinguishing subword for the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
Case 2.1.2. $u_{\operatorname{zeros}}$ is periodic.
The word $u_{\operatorname{zeros}}$ is of the following form: $u_{\operatorname{zeros}}=(1^{\alpha_{1}}01^{\alpha_{2}}0\cdots 1^{\alpha_{t}}0)^{\frac{l}{t}}$, where $\frac{l}{t}\geqslant 2$ is an integer and $t$ is minimal. Consider the word $u_{\operatorname{zeros}}^{\prime}=(1^{\alpha_{1}}01^{\alpha_{2}}0\cdots 1^{\alpha_{t}}0)(10)^{l-t}$. The word $1^{\alpha_{1}}01^{\alpha_{2}}0\cdots 1^{\alpha_{t}}0$ contains $\frac{n_{1}}{\frac{l}{t}}$ occurrences of 1, hence $u_{\operatorname{zeros}}^{\prime}$ contains $\frac{n_{1}\cdot t}{l}+l-t$ occurrences of 1 and $l$ occurrences of 0. The function $f(t)=\frac{n_{1}\cdot t}{l}+l-t$ is increasing and $t\leqslant\frac{l}{2}$ (since $\frac{l}{t}\geqslant 2$ is integer), so $u_{\operatorname{zeros}}^{\prime}$ contains at most $\frac{n_{1}+l}{2}$ occurrences of 1. We add to $u_{\operatorname{zeros}}^{\prime}$ all blocks $0^{x}$ from $u_{\operatorname{fix}}$ defined in the beginning of the proof. Let $u_{\operatorname{zeros}}^{\prime\prime}$ denote the obtained subword of $u$. There is only one way to take $u_{\operatorname{zeros}}^{\prime\prime}$ in $u$ modulo the selection of 1’s and 0’s from blocks from which we do not take all symbols. Since $l\leqslant n_{\operatorname{short}}-y+a+1$, we have $n=n_{1}+ax+n_{\operatorname{short}}\geqslant n_{1}+ax+l+y-a-1$. From these inequalities and (3) it follows that
$|u_{\operatorname{zeros}}^{\prime\prime}|+y+1\leqslant\frac{n_{1}+l}{2}+l+\left(\frac{a+2}{2}\right)(x-1)+y+1=$
$\frac{n_{1}+ax+l+y-a-1}{2}+l+x+\frac{y}{2}+\frac{1}{2}<\frac{n}{2}+\frac{n}{4}-\frac{ax-2x-2a+4}{2}+\frac{3x}{2}=$
@@
since $a\geqslant 5$ and $x\geqslant 2$. Applying Proposition 3.11 to the word $u_{\operatorname{zeros}}^{\prime\prime}$, we get that there exists a distinguishing subword of length at most $\frac{3}{4}n+4$ for $u$ and $v$.
Case 2.2. $a=4$.
In this case the proof is similar to the proof in Case 2.1. All inequalities from Case 2.1 hold true, except for the inequality (5) in Case 2.1.2. If we prove that $l\leqslant\frac{n}{4}-\frac{ax-2x-2a+4}{2}-\frac{y}{2}+1$, then we can rewrite inequality (5) as
$|u_{\operatorname{zeros}}^{\prime\prime}|+y+1\leqslant\frac{n_{1}+ax+l+y-a-1}{2}+l+x+\frac{y}{2}+\frac{1}{2}<\frac{n}{2}+\frac{n}{4}-\frac{ax-2x-2a+4}{2}$-
$\frac{y}{2}+1+x+\frac{x}{2}=\frac{3}{4}n-\frac{ax-2a-4x+2}{2}=\frac{3}{4}n-\frac{(a-4)(x-2)}{2}+3=\frac{3}{4}n+3,$
which gives us the proof in this case similarly to the proof in Case 2.1.
@@
$l\leqslant\frac{n}{4}-x+3-\frac{y}{2}.$ ($3^{\prime}$)
We now consider several cases according to the number of blocks of 0’s of length at least $\frac{y}{2}$ in $u$ and in $v$. Since in Case 2.2, the one we consider now, we have $a=4$, i.e., the number of blocks of 0’s of length $x$ (the longest blocks) in $u$ is $4$, $u_{\operatorname{long}}$ is not special by the conditions of the lemma, we have that the number of blocks of length at least $\frac{y}{2}$ in $u$ is at least $5$.
If there are at least six blocks in $u$ with lengths at least $0^{\frac{y}{2}}$, i.e. four blocks $0^{x}$, one block $0^{y}$ and at least one block $0^{\frac{y}{2}}$, then we can rewrite inequality (3) as $l\leqslant n_{\operatorname{short}}-y-(\frac{y}{2}-1)+1+a$. So, in this case we have $l\leqslant\frac{n}{4}-\frac{ax-2x-2a+4}{2}-\frac{y}{2}+1$.